OpenStudy (anonymous):

What is the total surface area of a cylinder if its radius is x + 2 and its height is x + 5?

OpenStudy (anonymous):

if it is a closed cylinder ,then surface area=area of curved surface+circular area of top+circular area of bottom.

OpenStudy (anonymous):



OpenStudy (anonymous):

I know how to find the area of a cylinder, I just don't know how to do it with binomials :(

OpenStudy (anonymous):

\[curved~surface=2\times \pi \left( x+2 \right)\left( x+5 \right)\] simplify it.

OpenStudy (anonymous):

6.28(x^2 +7) ?

OpenStudy (anonymous):

\[=2\pi \left( x^2+2x+5x+10 \right)=2\pi \left( x^2+7x+10 \right)\]

OpenStudy (anonymous):

ooohhhhhhhh ok! that's the surface area of the curved surface? Or the whole thing?

OpenStudy (anonymous):

surface area of base and top=2*(pi r^2) \[=2 \pi \left( x+2 \right)^2\] now find total surface area by adding the two .

OpenStudy (anonymous):

so 2*3.14(x^2 + 7x + 10) + 2*3.14(x + 2)^2

OpenStudy (anonymous):

4*3.14^2(x^3 + 7x + 12)